OpenStudy (anonymous):

Need Help w/ 3 Questions. [MEDAL & FAN] (Questions Inside)

OpenStudy (yoloshroom):

68 degrees for the first one :) 17x+16x+12x=180 45x=180 x=4 17(4)=68

OpenStudy (anonymous):

Ok thanks for adding the steps w/ the answer! What about the next one?

OpenStudy (yoloshroom):

(2x+7)/(3x-4)=11/2 Multiply both sides by the LCD : 2(3x-4) so you get.. 4x+14=33-44 Move X to one side, so -33x+4x=-44-14 combine like terms.. -29x=-58 divide by 29... -x=-2 Multiply by negative 1 to make X positive... x=2

OpenStudy (yoloshroom):

lastly, (6x+1)/7=(18x-2)/14 we can divide (18-2)/14 by 2 because all terms are divisible by two so (6x+1)/7=(9x-1)/7 mulitply both sides by 7 to remove the fractions. 6x+1=9x-1 move x to one side. 6x-9x=-1-1 combine like terms -3x=-2 divide by -3 x=-2/-3 x=-2/-3 negative/negative = positive so, x=2/3